Design a pair of straight gears, specifying the module the number of teeth, the pitch diameters, the distance between centers, the face width and the material with which the gears should be manufactured, if they will be used as part of a transmission to move A mobile sand sieve. The wheel should rotate at 450 rpm +/- 5%, having as an energy source an internal combustion engine that rotates at 2500 rpm with 80 H.P. The distance between centers should not be more than 200 mm. The transmission is expected to work for 5 continuous years without maintenance stops, in 10 hours / day, with 99.9% reliability. Evaluate your design with the criteria of resistance to bending and stinging. Specify sprocket and wheel construction materials. JUSTIFY ALL DESIGN DECISIONS.
Answer #1

THIS ANSWER IS TO EVALUATE THE DESIGN WITH THE CRITERIA OF RESISTANCE TO BENDING AND STINGING.450 I The fiven data Speed 5=4500pm.1/50.. E Engine Speed - 2500rpm; so H.P * Distance Between Centers more than 200mm VelociVelocity factor Cu = 3/879.42 +0.3.414 . 99-0.475 -0.50/400 4m YG = 0.175-000210 JOG=14ompa= 190 N/mm2 9252.67 = (140X0.24) r